OpenStudy (anonymous):

A culture started with 3,000 bacteria. After 6 hours, it grew to 3,300 bacteria. Predict how many bacteria will be present after 13 hours. Round your answer to the nearest whole number

OpenStudy (anonymous):

HELP?

OpenStudy (anonymous):

\[\frac{3300}{3000}=1.1\] so it increases by \(10\%\) every 6 hours compute \[\large 3000\times (1.1)^{\frac{13}{6}}\]

OpenStudy (anonymous):

i get \(3688\) http://www.wolframalpha.com/input/?i=3000%281.1%29^%2813%2F6%29

OpenStudy (anonymous):

I don't think that's right. Seems to me the bacteria will be growing exponentially.

OpenStudy (anonymous):

of course it is growing exponentially that is why \(\frac{13}{6}\) is in the exponent !